# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Unknown Dynamics

### v1

@@ Abstract
Safe maneuvering capability is critical for mobile robots in complex environments. However, robotic system dynamics are often time-varying, uncertain, or even unknown during the motion planning and control process. Therefore, many existing model-based reinforcement learning (RL) methods could not achieve satisfactory reliability in guaranteeing safety. To address this challenge, we propose a two-level Vector Field-guided Learning Predictive Control (VF-LPC) approach that guarantees safe maneuverability. The first level, the guiding level, generates safe desired trajectories using the designed kinodynamic guiding vector field, enabling safe motion in obstacle-dense environments. The second level, the Integrated Motion Planning and Control (IMPC) level, first uses the deep Koopman operator to learn a nominal dynamics model offline and then updates the model uncertainties online using sparse Gaussian processes (GPs). The learned dynamics and game-based safe barrier function are then incorporated into the learning predictive control framework to generate near-optimal control sequences. We conducted tests to compare the performance of VF-LPC with existing advanced planning methods in an obstacle-dense environment. The simulation results show that it can generate feasible trajectories quickly. Then, VF-LPC is evaluated against motion planning methods that employ model predictive control (MPC) and RL in high-fidelity CarSim software. The results show that VF-LPC outperforms them under metrics of completion time, route length, and average solution time. We also carried out path-tracking control tests on a racing road to validate the model uncertainties learning capability. Finally, we conducted real-world experiments on a Hongqi E-HS3 vehicle, further validating the VF-LPC approach’s effectiveness.
@@ I Introduction
For motion planning and tracking control tasks, prior information or desired paths are often required to guide robots’ motion. As an effective and efficient method, guiding vector field (GVF) techniques have been recently studied to realize path-following or obstacle-avoidance tasks successfully for robots like fixed-wing airplanes[1], miniature air vehicles[2], quadcopters[3], etc. The guidance feature of GVF manifests as follows[4, 5]: i) simple kinematic models of robots are considered (such as single or double integrator models [6, 7]); ii) GVF provides guidance signals to be tracked by the inner-loop dynamic controller (see Fig. 1). Therefore, the guidance level (i.e., the design of the GVF) and the control level can be separately designed. Specifically, GVF is designed to converge its integral curves to the desired geometric paths kinematically, and an effective robot dynamic controller is derived by exploiting the convergence properties of the integral curves under consideration of the robot’s dynamics. Therefore, obtaining sufficiently fast and precise dynamic controllers is important, as they must effectively respond to the kinematic controls (such as linear and angular velocities provided by GVFs). On the other hand, the dynamic controllers must take into account safety constraints. Considering the two aspects, it is imperative to propose an Integrated Motion Planning and Control (IMPC) approach. An IMPC approach generally considers the safety constraints and dynamics control, and generates controls directly transmitted to the low-level execution level. However, designing adaptive IMPC approaches for real-world robots with unknown dynamics is challenging, particularly in obstacle-dense scenarios.
@@
Figure 1: The proposed control architecture. In this figure, the GVF-based path-following controller generates desired linear and angular velocities $v$ and $\omega$; $F$ and $M$ are forces and moments provided by the IMPC module.
@@
Recent works on realizing reinforcement learning (RL) based IMPC were widely studied[8, 9, 10]. RL-based IMPC algorithms generally design proper reward functions based on priori information or desired paths to learn optimal policies. With the guidance information, existing endeavors primarily realize safe tracking control by designing reward functions encompassing state errors, control inputs, and safety terms. However, most of the studies assume that the desired paths are provided in advance, which is not valid when robots operate in different unstructured scenarios. When the desired path is constituted by a straight line from the starting point to the endpoint, it may traverse many obstacles in obstacle-dense environments. This increases the complexity of designing RL-based IMPC algorithms, posing challenges to achieving near-optimal performance. Therefore, providing guidance for RL methods is very important. Despite the above challenges, current RL-based IMPC has shown effectiveness and efficiency for robots with nonlinear system dynamics[8, 9, 10]. Letting RL-based IMPC approaches work as the IMPC structure in Fig. 1 is still promising if the above-discussed challenges can be well addressed. Namely, a GVF provides guidance for RL-based IMPC approaches.
@@
There exist two categories of RL-based IMPC studies; i.e., model-based and model-free ones. Obtaining a precise dynamic model is nontrivial due to internal factors such as system nonlinearity, and external factors such as uneven terrain, and slippery surfaces. The advantage of model-free RL methods lies in their independence from precise models, but they require many training samples for learning, resulting in much training time and very high risk. Efficient model-based RL approaches have shown their effectiveness in real-world applications[9, 10]. However, since the real-world robot dynamics are often uncertain or even unknown, model-based RL (MBRL) approaches struggle to achieve satisfactory reliability. This is due to: 1) The accuracy and adaptability of current data-driven modeling methods are insufficient. 2) The online learning efficiency of MBRL is too low to achieve fast convergence.
@@
To realize motion planning of mobile robots with unknown dynamics in obstacle-dense scenarios, we propose a Vector Field-guided Learning Predictive Control (VF-LPC) approach. Specifically, the guiding level plans preliminary trajectories fast to avoid dense static obstacles. The optimal performance under the unknown system dynamics and the safety constraints from (suddenly appearing) moving obstacles are optimized by solving the Hamilton–Jacobi–Bellman (HJB) equation online in prediction horizons. In particular, we introduce a sparsification technique in the model compensation and finite-horizon actor-critic learning processes to improve the online efficiency and control performance. The contributions of this paper are summarized as follows:
@@
    By adding virtual obstacles, the modified and improved discrete-time composite vector field adopted by our VF-LPC approach can satisfy robot kinodynamic constraints. In addition, the modified vector field does not suffer from the deadlock problem, which usually exists in traditional composite vector fields. Moreover, VF-LPC can deal with (suddenly appearing) moving obstacles by introducing a game-based barrier function.
@@
    The VF-LPC approach can update online the uncertain dynamics of a fully data-driven model trained by the deep Koopman operator. It reduces the differences between the real and learned system dynamics models when the environment is time-varying or the system dynamics are learned inaccurately. Therefore, it improves the control performance and guarantees safety.
@@
    The VF-LPC approach can achieve near-optimal safety control for robots with unknown dynamics. The approach achieves higher computational efficiency and obtains more reliable solutions than advanced model predictive control (MPC) methods in solving nonlinear optimization problems.
@@
    Our proposed VF-LPC approach not only has theoretical guarantees but also has been demonstrated effective in practice since it has been validated by extensive simulations and experiments with a Hongqi E-HS3 vehicle in an environment with moving obstacles.
@@ II Related Work
We first discuss several modeling methods and then present a literature review on popular MPC and MBRL approaches for the safety control of systems with uncertain dynamics.
@@
Data-driven modeling. Current advanced data-driven modeling methods include least-squares[11], recurrent neural networks (NNs)[12], multi-layer perception (MLP)[13], neural networks[14], etc. As a linear operator, the Koopman operator-based modeling methods[15, 16, 17, 18] can establish linear time-invariant system dynamics. Impressed by such a property, dynamic mode decomposition (DMD)[15], extended DMD (EDMD)[16], and kernel-based DMD[17] have received much attention in recent years. However, the modeling performance of the Koopman operator relies heavily on the observable function design. Consequently, approaches using NNs for automated observable function construction [19, 20, 21] were proposed to improve the modeling accuracy and have been validated to be effective through numerical simulations. To improve the modeling accuracy, Xiao et al. [22] proposed a deep Koopman operator method for identifying linear time-invariant vehicle dynamic models to track reference velocities. Additionally, vehicle control experiments were conducted in [23] to validate the effectiveness of the deep direct Koopman (DDK) method. Unlike these works, we further consider improving the online adaptability of offline-trained dynamics models by learning the uncertain dynamics of the offline-trained Koopman model.
@@
Sampling-based and optimization-based motion planning algorithms for robots with uncertain dynamics. These algorithms necessitate the generation of paths/trajectories that conform to dynamic feasibility, collision avoidance, and uncertain dynamics. Lars et al. proposed a chance-constrained path planning method[24], which employs a probabilistic representation of uncertainty, thereby constraining the probability of collisions or deviations from specified regions within threshold limits. Several studies have integrated chance constraints with sampling-based Rapidly-exploring Random Trees (CC-RRT) methods, presenting efficient path planning capability in densely cluttered obstacle environments[25]. Gaussian processes (GPs) were employed for determining dynamically feasible paths and CC-RRT for establishing probabilistically feasible paths [26]. In [27], Kalman filtering was utilized for state estimation and risk-aware safety constraints arising from estimation errors were introduced into stochastic optimal control problems. However, these approaches fail to guarantee optimality due to the lack of consideration of the robot’s motion dynamics constraints. Utilizing model predictive control (MPC) to address motion planning under robot dynamic uncertainties typically involves objective functions incorporating Conditional Value-at-Risk (CVaR) measures[28, 29]. In summary, under unknown system dynamics, it is still challenging to efficiently realize safe and non-conservative control.
MBRL for motion planning of robots with uncertain dynamics. Some offline MBRL approaches complete complex tasks by using enormous training samples. Diehl et. al. tried to jointly solve multi-level problems of automated driving[30] in a simulator using offline collected real-world datasets. In [31], the new AEMCARL framework enables collision-free robot navigation in complex scenarios through adaptive modeling and learning. Besides these tasks, relevant efforts focus on specific tracking-avoiding tasks. In [9], our previous work devised an efficient online batch-mode RL approach to realize motion planning of uncertain vehicle dynamics. Zhang et al. [10] proposed receding-horizon RL with NN-based model compensation for realizing motion planning of intelligent vehicles. In this work, we focus on robot motion-planning tasks and design a new framework that further considers the unknown dynamics in obstacle-dense environments for realizing kinodynamic motion planning and control.
@@ III-A Composite Vector Field
As shown in Fig. 2, the elements of the composite vector field are illustrated in detail. A reference path $\mathcal{P}$ is provided initially and may be occluded by obstacles, and it is defined by
@@ III-D1 Composite Vector Field with Kinodynamic Constraints
The term “kinodynamic constraint” refers to the requirement that a robot will not collide with obstacles at different speeds. To address this issue, we have transformed it into the fulfillment of the maximum lateral acceleration.
@@ III-D2 Optimal Motion Planning to Avoid Moving Obstacles
Given the offline learned system (4), it is feasible to use it to design optimal controllers. However, the interaction environments may be time-varying, causing the system dynamics to be uncertain. We can rewrite the exact system dynamics as a data-driven Koopman model adding an uncertain part by
@@
where $0<\gamma\leq 1$, $L(\tilde{x}_{\tau},\tilde{u}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\tilde{u}_{\tau}^{\top}R\tilde{u}_{\tau}$ is the cost function, $Q\succeq 0\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is positive semi-definite, and $R\succ 0\in\mathbb{R}^{n_{u}\times n_{u}}$ is positive definite. The second subproblem is how the robot can avoid moving obstacles. Combining the above two subproblems, we formulate the optimal motion planning problem (OMP) as below:
@@
(OMP) Design an optimal safe controller for the robot with uncertain system dynamics such that it
@@
IV Vector Field Guided Receding Horizon Reinforcement Learning for Mobile Robots with Unknown System Dynamics
It is essential to generate local collision-free trajectories for guiding robots’ motion in obstacle-dense scenarios, which could improve safety and simplify the design of RL dynamic controllers. Motivated by this aspect, we design a guiding vector field that considers dynamic constraints and excludes the deadlock problem (i.e., singular points). This part is illustrated by the safety guiding module in Fig. 4. Considering safety when robots track the desired trajectory, we must deal with the movements of (suddenly appearing) moving obstacles and the uncertainties of the nominal deep Koopman model. To this end, we develop an online receding-horizon reinforcement learning (RHRL) approach that employs a game-based exponential barrier function and a fast model compensation scheme. This part is illustrated by the learning predictive control module of Fig. 4. The details of each module and its sub-modules will be illustrated in the following subsections.

### v2

# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Uncertain Dynamics

@@ Abstract
In obstacle-dense scenarios, providing safe guidance for mobile robots is critical to improve the safe maneuvering capability. However, the guidance provided by standard guiding vector fields (GVFs) may limit the motion capability due to the improper curvature of the integral curve when traversing obstacles. On the other hand, robotic system dynamics are often time-varying, uncertain, and even unknown during the motion planning process. Therefore, many existing kinodynamic motion planning methods could not achieve satisfactory reliability in guaranteeing safety. To address these challenges, we propose a two-level Vector Field-guided Learning Predictive Control (VF-LPC) approach that improves safe maneuverability. The first level, the guiding level, generates safe desired trajectories using the designed kinodynamic GVF, enabling safe motion in obstacle-dense environments. The second level, the Integrated Motion Planning and Control (IMPC) level, first uses a deep Koopman operator to learn a nominal dynamics model offline and then updates the model uncertainties online using sparse Gaussian processes (GPs). The learned dynamics and a game-based safe barrier function are then incorporated into the LPC framework to generate near-optimal planning solutions. Extensive simulations and real-world experiments were conducted on quadrotor unmanned aerial vehicles and unmanned ground vehicles, demonstrating that VF-LPC enables robots to maneuver safely.
@@ I Introduction
Prior information or desired paths are often required to guide robots’ motion. As an effective and efficient method, guiding vector field (GVF) techniques have been recently studied to realize path-following or obstacle-avoidance tasks successfully for robots like fixed-wing airplanes[1], unicycle-type vehicles[2], unmanned aerial vehicles[3], etc. The GVF typically addresses motion planning and control tasks in the following manner[4]: (1) simple kinematic models of robots are considered (such as single or double integrator models); (2) GVF provides guidance signals to be tracked by the inner-loop dynamic controller). Therefore, the guidance level (i.e., the design of the GVF) and the control level can be separately designed. However, the proposed GVFs such as[5] may generate guidance signals with drastic variations when crossing static obstacles, posing a challenge to the robot’s maneuvering capabilities. On the other hand, robot dynamics and safety constraints (e.g., avoidance of suddenly appearing obstacles) are critical in improving the robots’ maneuvering capabilities. When considering the robotic dynamics constraint separately at planning and control levels, the issues of consistency and optimality have to be carefully addressed for real-world applications. Therefore, it is crucial to introduce an Integrated Motion Planning and Control (IMPC) approach to generate solutions transmitted to robot dynamics directly. In light of the above two aspects, designing adaptive IMPC approaches that leverage the improved GVF as guidance for real-world robots with uncertain/unknown dynamics is promising, particularly in obstacle-dense scenarios; An illustration diagram is presented in Fig. 1.
@@
Figure 1: The proposed integrated motion planning and control architecture. In this figure, the guidance module generates desired linear and angular velocities $v$ and $\omega$; $F$ and $M$ are forces and moments used to control a robot directly.
@@
Optimization-based motion planning methods have recently been studied for realizing IMPC [6, 7]. It is crucial to estimate model uncertainties in real-time to address the challenge of uncertain robotic dynamics impacting model predictive control (MPC) methods. This can be achieved through online estimation techniques, while solving nonlinear or non-convex optimization problems online may pose reliability and computational intensity issues. Reinforcement learning (RL) and adaptive dynamic programming (ADP) are promising in solving optimal planning and control problems[8],[9],[10]. Recent works on realizing RL-based IMPC have been studied[9, 10]. They generally design proper reward functions based on priori information or desired paths to learn optimal policies. With the desired paths, existing endeavors primarily realize safe tracking control by designing reward functions encompassing state errors, control inputs, and safety terms. However, in obstacle-dense scenarios, obtaining the desired safe paths is not easy. When the desired path is constituted by a straight line from the starting point to the endpoint, it may traverse many obstacles in obstacle-dense environments. Furthermore, the desired speed or angular velocity for these paths needs to be computed separately. The two aspects increase the complexity of designing RL-based IMPC algorithms, posing challenges to achieving near-optimal performance. Therefore, providing guidance for them is very important. Despite the above difficulties, current RL-based IMPC has shown effectiveness and efficiency for robots with nonlinear system dynamics[8],[9],[10]. Letting RL-based IMPC approaches work as the IMPC structure in Fig. 1 is still promising if the above-discussed challenges can be well addressed. Namely, a GVF provides kinematic guidance (e.g. linear and angular velocities) for RL-based IMPC approaches.
@@
There exist two main categories of RL-based IMPC studies; i.e., model-based and model-free ones. Obtaining a precise dynamic model is nontrivial due to internal factors such as system nonlinearity, and external factors such as uneven terrain, and slippery surfaces. The advantage of model-free RL methods lies in their independence from precise models. However, they still lack generalization ability to the unseen scenarios and data-efficiency of policy learning. Efficient model-based RL (MBRL) approaches have shown their effectiveness in real-world applications[9, 10]. Since real-world robot dynamics are often uncertain and even unknown, MBRL approaches struggle to achieve satisfactory reliability. This is due to: (1) The online adaptability of current data-driven modeling methods is insufficient; (2) The safety terms in the reward function may lead to policy divergence of RL-based motion planning algorithms. Motivated by the two challenges, we propose compensating for the data-driven model online and adopting a receding-horizon actor-critic framework (also called Learning Predictive Control[11], LPC) to ensure the convergence of planning policies in the prediction horizons.
@@
To realize motion planning of mobile robots with uncertain/unknown dynamics in obstacle-dense scenarios, we propose a Vector Field-guided Learning Predictive Control (VF-LPC) approach. Specifically, the guiding level plans preliminary trajectories fast to avoid dense static obstacles. The optimal performance under the uncertain/unknown system dynamics and the safety constraints from (suddenly appearing) moving obstacles are optimized by solving the Hamilton–Jacobi–Bellman (HJB) equation online in prediction horizons. In particular, we introduce a sparsification technique in the model compensation and finite-horizon actor-critic learning processes to improve online efficiency and IMPC performance. The contributions of this paper are summarized as follows:
@@
    The VF-LPC approach can achieve near-optimal motion planning for mobile robots with uncertain/unknown dynamics in obstacle-dense environments. The approach achieves higher computational efficiency and obtains more reliable solutions than advanced model predictive control (MPC) and RL methods in solving nonlinear optimization problems (see Sections IV-D and VI).
@@
    Our proposed VF-LPC approach not only has theoretical guarantees but also has been demonstrated effective in practice since it has been validated by extensive simulations and experiments with quadrotor unmanned aerial vehicles (UAVs) and a Hongqi E-HS3 vehicle (see Sections V and VI).
@@
    The VF-LPC approach can update online the uncertain dynamics of a fully data-driven model trained by the deep Koopman operator. It reduces the differences between the real and learned system dynamics models when the environment is time-varying or the system dynamics are learned inaccurately. Therefore, it improves the planning performance and guarantees safety (see Section IV-B).
@@
    By adding virtual obstacles, the modified and improved discrete-time composite vector field adopted by our VF-LPC approach can satisfy robot kinodynamic constraints. In addition, the modified vector field does not suffer from the deadlock problem, which usually exists in traditional composite vector fields. Moreover, VF-LPC can deal with (suddenly appearing) moving obstacles by introducing a game-based barrier function (see Sections IV-A, IV-C).
@@ II Related Work
We first discuss several modeling methods and then present a literature review on MPC and learning-based approaches for motion planning of robots with uncertain dynamics.
@@
Data-driven modeling. Current advanced data-driven modeling methods include least-squares[12], recurrent neural networks (NNs)[13], multi-layer perception (MLP)[14], neural networks[15], etc. As a linear operator, the Koopman operator-based modeling methods[16],[17],[18] can establish linear time-invariant system dynamics. Impressed by such a property, dynamic mode decomposition (DMD)[16], extended DMD (EDMD)[17], and kernel-based DMD[18] have received much attention in recent years. However, the modeling performance of the Koopman operator relies heavily on the observable function design. Consequently, approaches using NNs for automated observable function construction [19, 20] were proposed and have been validated to be effective through numerical simulations. To improve the modeling accuracy, Xiao et al. [21] proposed a deep direct Koopman (DDK) method for identifying linear time-invariant vehicle dynamic models. Unlike these, we further consider improving the online adaptability of offline-trained dynamics models by learning the uncertain dynamics of the offline-trained Koopman model.
@@
Sampling-based and optimization-based motion planning algorithms for robots with uncertain dynamics. Several studies have integrated chance constraints with sampling-based Rapidly-exploring Random Trees (CC-RRT) methods, presenting efficient path planning capability in densely cluttered obstacle environments [22, 23]. Gaussian processes (GPs) were employed for determining dynamically feasible paths and CC-RRT for establishing probabilistically feasible paths[23]. These approaches fail to guarantee optimality due to the lack of consideration of the robot’s motion dynamics.
Under uncertain dynamics, optimization-based motion planning methods typically involve objective functions incorporating Conditional Value-at-Risk (CVaR) measures[24],[25]. Nakka et. al. handle the motion planning problem of chance-constrained nonlinear stochastic systems by deriving a surrogate problem with convex constraints[6]. Zhu et. al. designed a chance-constrained nonlinear MPC method to solve collision avoidance problems of multi-robots under various uncertainties like motion disturbance[26]. To realize real-time optimization, they developed a tight bound for the approximation of collision probability. In[27], Lew et. al. proposed a robust trajectory optimization method for nonlinear systems with model uncertainty and disturbances. Especially, it is capable of processing non-convex obstacle constraints. In [28], Kalman filtering was utilized for state estimation, and risk-aware safety constraints arising from estimation errors were introduced into stochastic optimal motion planning problems. Many of the above studies process various uncertainties like sensing, obstacle motion, robot dynamics, etc. According to [29], two major ideas are considered in the area of feedback motion planning, i.e., probabilistic guarantees on safety and bounded models of uncertainties. In this paper, the unknown system dynamics are modeled with the previous deep Koopman operator[21]. We consider the bounded uncertainty of data-driven system dynamics and compensate online for it.
Learning-based motion planning for robots with uncertain dynamics. As discussed in[30], RL-based approaches to addressing obstacle avoidance problems under uncertainties typically fall into two categories: one involves the endeavor to construct stochastic models of the uncertain dynamics inherent in robotic systems, leveraging the resultant probabilistic models for planning or policy learning; the other entails devising plans that account for worst-case scenarios. Regarding the first category, a Gaussian process model was used in the policy search framework of PILCO, thereby capturing the system dynamics and estimating the probability of safe constraints violation[31]. During the policy learning process, the candidate policies are optimized toward the safer directions with low risks. In [32], the Monte Carlo motion planning method was proposed to sample feasible trajectories under uncertainties, thereby fulfilling probabilistic collision avoidance constraints. Model-based motion planning under uncertain dynamics can be found in [9],[10]. Regarding the second category, Snyder et. al. proposed a trust-region-based online learning algorithm with provable regret bounds by minimizing worst-case regret[30]. To realize kinodynamic motion planning and control, we design an IMPC framework leveraging the guidance from the vector field and further consider the uncertainties of the data-driven model in obstacle-dense environments.
@@ III-A Composite Vector Field
In Fig. 2, the elements of the composite vector field are illustrated in detail. A reference path $\mathcal{P}$ is provided initially and may be occluded by obstacles, and it is defined by
@@ III-D1 Composite Vector Field with Kinodynamic Constraints
[…]
@@ III-D2 Optimal Motion Planning to Avoid Moving Obstacles
Given the offline learned system (4), it is feasible to use it to design optimal IMPC. However, the interaction environments may be time-varying, causing the system dynamics to be uncertain. We can rewrite the exact system dynamics as a data-driven Koopman model adding an uncertain part by
@@
where $0<\gamma\leq 1$, $L(\tilde{x}_{\tau},\tilde{u}_{\tau})=\tilde{x}_{\tau}^{\top}Q\tilde{x}_{\tau}+\tilde{u}_{\tau}^{\top}R\tilde{u}_{\tau}$ is the cost function, $Q\succeq 0\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is positive semi-definite, and $R\succ 0\in\mathbb{R}^{n_{u}\times n_{u}}$ is positive definite. The second subproblem is how the robot can avoid moving obstacles. Combining the two subproblems, we formulate the optimal OMP as below:
@@
(OMP) Design an optimal IMPC for the robot with uncertain system dynamics such that it
@@
IV Vector Field Guided Receding Horizon Reinforcement Learning for Mobile Robots with Uncertain System Dynamics
It is essential to generate local collision-free trajectories for guiding robots’ motion in obstacle-dense scenarios, which could improve safety and simplify the design of RL-based IMPC. Motivated by this aspect, we design a guiding vector field that considers dynamic constraints and excludes the deadlock problem (i.e., singular points). This part is illustrated by the safety guiding module in Fig. 4. Considering safety when robots track the desired trajectory, we must deal with the movements of (suddenly appearing) moving obstacles and the uncertainties of the nominal deep Koopman model. To this end, we develop an online receding-horizon reinforcement learning (RHRL) approach that employs a game-based exponential barrier function and a fast model compensation scheme. This part is illustrated by the learning predictive control module of Fig. 4. The details of each module and its sub-modules will be illustrated in the following subsections.